OpenStudy (anonymous):

How many yards are in 2 ½ miles if there are 5280 feet in one mile?

OpenStudy (king):

1 mile =5280 feet we know that, 1 yard=3 feet So, 1mile =5280/3 1 mile=1760 yards so, 2.5 miles=1760*2.5 =4400 yards